Previous

Calendar

July

2017

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

25

JD176

26

JD177

27

JD178
GLEES Site Visit (JF): Details ,

both quad-disk ports are still co-located side-by-side at 22.65m

swapped which quad-disk is connected to which transducer

Gray Box (SN 9301150) connected to clear quad-disk: changed from tilt=-9 deg to tilt=0 deg

Yellow Box (SN 950329) connected to Qualmetrics (opaque) quad-disk: tilt=0 deg (no change)

NOTE: the pressure ports didn't move; the Qualmetrics quad-disk is located closer to the sonic; the older clear pressure head is mounted further from the sonic and lateral to the north

28

JD179

29

JD180

30

JD181
Site Visit (SpB): Details , Photos (36)

Span cylinder at xxx psi. N2 at 300 ---> 1950 psi

SpB Jeep

Misty in AM, Partly-cloudy in PM

1

JD182

2

JD183

3

JD184
SpB Off

4

JD185
Independence Day
GLEES Site Visit (JF): Details ,

disconnected the 22.65m co-located side-by-side quad-disk ports from BOTH transducers (tubing is still intact)

5

JD186

6

JD187
SpB Returns

7

JD188

8

JD189

9

JD190
at 11 PM, urquell auto-reboot.

(Full "Buck" Moon)

10

JD191
Start of the NSF Fluxcourse at MRS.

11

JD192

12

JD193
Site Visit (SpB, DB, M, RM, DM, JK, KN, +): Details , Photos (24)

Span cylinder at xxx psi. N2 at xxx psi

at XX:XX MST, changed mV output range for UofU PAR sensors (rad_05.csi)

Flux class (site visit)

SpB Jeep

Overcast in AM, light rain in PM

13

JD194
MMM Retreat (SpB)

14

JD195
MMM Retreat (SpB)

15

JD196

16

JD197

17

JD198

18

JD199

19

JD200

20

JD201

21

JD202
End of the NSF Fluxcourse at MRS.

22

JD203

23

JD204

24

JD205

25

JD206

26

JD207

27

JD208

28

JD209

29

JD210

30

JD211

31

JD212

1

JD213

2

JD214

3

JD215

4

JD216

5

JD217